compute algebraicially the resultant of the following coplanar forces: 100 N at 30 degrees, 141.4N at 45 degrees and 100N at 240 degrees



1
Expert's answer
2021-11-04T10:15:13-0400

"x=100\cdot\cos30°+141.4\cdot\cos45°+100\cdot\cos240°="


"=136.59\ (N)"


"y=100\cdot\sin30°+141.4\cdot\sin45°+100\cdot\sin240°="


"=63.4\ (N)"


"F_{total}=\sqrt{136.59^2+63.4^2}=150.6\ (N)"